OpenStudy (anonymous):

What is the solution of this linear system? 4x+y=-1 y=-2x+3

OpenStudy (anonymous):

you can put equations on the same side (1) 4x+y=-1 (2) 2x+y=3 and then substract them (1)-(2) which is 2x=-4 which implies x=-2 and then you can find y . you could also put y=-2x+3 into the first equation resulting in 4x+(-2x+3)=-1 which gives 2x+3=-1 which is x=-2 and then you can get y by pluggin x into any of the equations

OpenStudy (anonymous):

is -1 and -2 the answer?

OpenStudy (anonymous):

because my multiple choice says -2, -7 or -2, 7

OpenStudy (anonymous):

if you put x=-2 in the equation y=-2x+3 you get y= -2*(-2)+3= 7 so the answer is -2,7 you forgot that (-2)*(-2)=4 a negative times a negative is positive